# On Fock covariance for product systems and the reduced Hao–Ng isomorphism problem by discrete actions

### v1

@@ 1.1. Fock covariance
Hilbertian representations play an important role in the study of algebraic structures and symmetries. One of the cornerstone examples is the Gel’fand–Raikov Theorem that “identifies” a locally compact group with its unitary representations. Further examples come from C*-dynamical systems, (topological) graphs, semigroups, and groupoids. Algebraic structures associated with discrete groups have been successfully unified under the theory of Fell bundles. Product systems over a unital subsemigroup $P$ of a discrete group $G$ provide a common context to describe irreversible transformations as a semigroup analogue of Fell bundles. There has been a great number of results in the area, perhaps too many to list here, and we will just comment on work as related to our results.
@@
Product systems were introduced by Arveson [2] disguised under duality, and later put in context by Dinh [10] for discrete subsemigroups of $\mathbb{R}_{+}$. Following Nica [42] and Pimsner [45], Fowler [18] proceeded to an in-depth study when $(G,P)$ is a quasi-lattice, that inspired a great number of subsequent works. It had been known that, unlike to the group case, taking plain isometric representations leads to intractable objects, e.g., the universal isometric C*-algebra of $\mathbb{N}\times\mathbb{N}$ is not even nuclear [40]. A richer structure is at hand when one considers the relations in the Fock representation of a quasi-lattice order, known as Nica covariance. In order to make full use of this additional feature, Fowler imposed several axioms on the product system resulting to a by-default Wick ordering. Kwaśniewski and Larsen [32, 33] extended this setup to the case of right LCM semigroups, and essentially this is the furthest one may go and still have a Wick ordering. Cuntz, Deninger and Laca [8] used the Fock model beyond Nica’s work [42] for $P=R\rtimes R^{\times}$, where Nica covariance may fail, but one can still use the Fock representation as the prototype for covariant relations.
@@
The coaction of the ambient group $G$ has been pivotal in the theory of product systems, in particular in relation to boundary quotients by Carslen, Larsen, Sims and Vittadello [7] and Sehnem [47, 48], as well as in semigroup algebras by Laca and Sehnem [34] and the first author, Katsoulis, Laca and Li [22]. Dor-On, the first author, Katsoulis, Laca and Li [11] noted that there is a canonical Fell bundle arising from the normal coaction on ${\mathcal{T}}_{\lambda}(X)$, denoted here by ${\mathcal{F}}{\mathcal{C}}X$. It appears that, while ${\mathcal{T}}_{\lambda}(X)$ is the reduced C*-algebra of ${\mathcal{F}}{\mathcal{C}}X$, it is not always the reduced C*-algebra for the universal C*-algebra ${\mathcal{T}}(X)$ of the representations of $X$ (as it happens for $P=\mathbb{Z}_{+}$). To make a distinction, we will write ${\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$ for the universal C*-algebra of ${\mathcal{F}}{\mathcal{C}}X$; a representation of $X$ that promotes to a *-representation of ${\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$ will be called Fock covariant. It was shown in [11] that Fock covariant representations coincide with Nica covariant representations for a compactly aligned product system $X$ over a right LCM semigroup $P$, and thus ${\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$ coincides with the universal Nica covariant C*-algebra ${\mathcal{N}}{\mathcal{T}}(X)$ in this case.
@@
Motivated by the Hao–Ng isomorphism problem, our first aim in this work is to provide a description of the equivariant Fock covariant injective representations of $X$. For our purposes, we will view a product system $X$ as a family of C*-correspondences $\{X_{p}\}_{p\in P}$ in a common ${\mathcal{B}}(H)$ that satisfies some natural semigroup relations, i.e., that $A:=X_{e}$ is a C*-algebra, and that
@@
This setup enables to still define left creation operators on the full Fock space ${\mathcal{F}}X$ giving rise to the Fock C*-algebra ${\mathcal{T}}_{\lambda}(X)$, while it recovers Fowler’s product systems in [18] which arise under the stronger saturation condition that $[X_{p}X_{q}]=X_{pq}$ for all $p,q\in P$. We note that most of the works for product systems assume non-degeneracy of the left action on $X$. Nevertheless, non-degeneracy is rarely used, although imposed by the saturation condition of Fowler [18]. The characterisation we provide is given in terms of the $\mathbf{K}_{{\mathbf{x}}}$-cores for ${\mathbf{x}}\in{\mathcal{J}}$. We then use the natural description of the fixed point algebra by an inductive limit of C*-algebras on the $\cap$-closed finite subsets of the constructible ideals of $P$ in ${\mathcal{J}}$, which are spanned by the $\mathbf{K}_{\bullet}$-cores.
@@
We square this characterisation with established results from the literature. Towards this end, we provide a direct connection with the equivariant Nica covariant injective representations when $P$ is a right LCM semigroup recovering the key result of [11], and with the equivariant Toeplitz semigroup representations of Laca and Sehnem [34] when $X$ arises from $P$ (note that injectivity is automatic in the latter case). As a consequence we get that Fock covariance coincides with Nica covariance for compactly aligned product systems over a right LCM semigroup, and with Toeplitz covariance for semigroup representations.
@@ 1.2. The Hao–Ng isomorphism problem
and then the C*-envelope theory of Katsoulis and Ramsey [27] can be implemented. In the absence of independence, it is unclear whether the *-isomorphism (1.5) still holds, but it does not exclude the possibility that the completely isometric isomorphism (1.6) is valid. A careful investigation of the arguments of [11, 12, 24, 25, 28] reveals that Fock covariance of the identity representation $X\rtimes_{\alpha,\lambda}{\mathfrak{H}}\hookrightarrow{\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ enables the completely isometric isomorphism (1.6). Our characterisation applies towards this resolution when ${\mathfrak{H}}$ is discrete without restrictions on the product system $X$ and the semigroup $P$.
@@ Acknowledgements.
The authors would like to thank Elias Katsoulis for his helpful comments and remarks on the draft of the manuscript. The second author acknowledges support from CaLISTA which funded the research visit of the second author to Newcastle University in October 2023. This manuscript will form part of the second author’s PhD thesis.
@@ 2.1. Coactions on C*-algebras
There is a direct connection between coactions on C*-algebras, gradings and Fell bundles. The reader is addressed to [16, 17, 46] where this theory is laid in full detail. Since the group $G$ is discrete, the coactions we consider are automatically non-degenerate in the sense of [46], see for example [11, Remark 3.2].
@@
together spanning a dense subset of ${\mathcal{C}}$, see the proof of [41, Proposition 2.6]. If, in addition, the map $({\operatorname{id}}\otimes\lambda)\circ\delta$ is faithful, then $\delta$ will be called normal. It follows that $\delta$ is normal if and only if ${\mathcal{C}}$ admits a reduced coaction by $G$, i.e., there is a faithful *-homomorphism $\delta_{\lambda}\colon{\mathcal{C}}\to{\mathcal{C}}\otimes\mathrm{C}^{*}_{\lambda}(G)$ such that
@@
If there exists a conditional expectation $E\colon{\mathcal{C}}\to{\mathcal{C}}_{e}$, then the subspaces are linearly independent, see [16, Theorem 3.3]. In this case the C*-grading is called topological. By definition, a coaction on a C*-algebra induces a topological C*-grading.
@@
for the canonical embedding. We use the same symbol $\Psi$ for the *-homomorphism of $\mathrm{C}^{*}_{\max}({\mathcal{B}})$ induced by a representation $\{\Psi_{g}\}_{g\in G}$ of ${\mathcal{B}}$ (as $\Psi\circ\hat{j}_{g}=\Psi_{g}$). By universality, we have that $\{\hat{j}_{g}\}_{g\in G}$ is an equivariant representation of ${\mathcal{B}}$, and in particular by [17, Proposition 17.9] we have that the map $\hat{j}_{g}\colon{\mathcal{B}}_{g}\to[\mathrm{C}^{*}_{\max}({\mathcal{B}})]_{g}$ is an isometric isomorphism. Hence, any Fell bundle arises as a C*-grading from some C*-algebra.
@@
and thus a normal coaction. By [16, Theorem 3.3] we have that, if $\Psi$ is an equivariant representation of ${\mathcal{B}}$ that is injective on ${\mathcal{B}}_{e}$, then there are equivariant canonical *-epimorphisms
@@
More generally, let $\Psi$ be an equivariant (possibly non-injective) representation of ${\mathcal{B}}$, then by the proof of [17, Proposition 21.4] (or by using the Fourier transform) we obtain that every $\Psi_{g}$ has closed range, and hence we have the induced Fell bundle
@@
commutative, see [17, Proposition 21.2, Proposition 21.3]. We will make use of the following folklore fact for Fell bundles.
@@
It is obvious that if the inclusion holds for all $g\in G$ then in particular it holds for $g=e$. Conversely, suppose that
@@
and so $x\in\ker\Psi_{2}$. Consequently we derive $x\in\ker\Psi_{2}\cap[\mathrm{C}^{*}_{\max}({\mathcal{B}})]_{g}$, as required.
@@
In that case, the canonical quotient map $q_{\mathcal{I}}$ is equivariant, i.e., the coaction $\delta$ descends to a coaction on ${\mathcal{C}}/{\mathcal{I}}$, see [16, Proposition 3.11] and [7, Proposition A.1] for the full details. The following proposition is perhaps folklore, and we include a proof for completeness.
@@
On the other hand, since $q_{\mathcal{I}}({\mathcal{B}})$ is a topological C*-grading of $\mathrm{C}^{*}_{\max}({\mathcal{B}})/{\mathcal{I}}$, by [17, Theorem 19.5] there exists a canonical *-epimorphism
@@ 2.2. Semigroups
We will require some elements on semigroup algebras and right ideals of a semigroup from the work of Li [36], while we also fix notation. For a unital discrete left-cancellative semigroup $P$, we let the left creation (isometric) operators given by
@@
By definition we have that $p^{-1}P=P$. We write ${\mathcal{J}}$ for the smallest family of right ideals of $P$ containing $P$ and $\emptyset$ that is closed under left multiplication, taking pre-images under left multiplication (as in the sense above) and finite intersections, i.e.,
@@
The right ideals of $P$ in ${\mathcal{J}}$ are called constructible. By the proof of [36, Lemma 3.3] we obtain a reduced form for the elements of ${\mathcal{J}}$, since
@@
Henceforth we will assume that $P$ is a unital subsemigroup of a discrete group $G$. In this case we have that
@@
and therefore inductively we obtain that
@@
is a reduced coaction, and thus it lifts to a normal coaction on $\mathrm{C}^{*}_{\lambda}(P)$. The induced faithful conditional expectation on $\mathrm{C}^{*}_{\lambda}(P)$ coincides with the compression to the diagonal $\ell^{\infty}(P)$. In the presence of an overlying group there is an explicit formula for the projection $E_{[{\mathbf{x}}]}$ for ${\mathbf{x}}\in{\mathcal{J}}$, obtained in [36, Lemma 3.1], i.e.,
@@ 2.3. C*-correspondences
The theory of Hilbert modules over C*-algebras is well-developed. The reader is addressed to [35, 39] for an excellent introduction to the subject.

### v2

@@ 1.1. Fock covariance
Hilbertian representations play an important role in the study of algebraic structures and symmetries. One of the cornerstone examples is the Gel’fand–Raikov Theorem that “identifies” a locally compact group with its unitary representations, while further examples arise naturally in the context of invertible transformations on a (possibly noncommutative) state space. In this endeavour, algebraic structures associated with discrete groups have been successfully unified under the theory of Fell bundles. Product systems over a unital subsemigroup $P$ of a discrete group $G$ provide a common context to describe irreversible transformations as a semigroup analogue of Fell bundles. Examples include subshift spaces, rank one or higher rank graphs, topological graphs, and semigroup actions on C*-algebras (in particular in relation to transfer operators), to mention but a few beyond the Fell bundles setup.
@@
Product systems were introduced by Arveson [2] disguised under duality, and later put in context by Dinh [12] for discrete subsemigroups of $\mathbb{R}_{+}$. Nica [46] considered operator algebras related to semigroups that admit a type of order structure, i.e., quasi-lattice semigroups, for which a representation by left translations was well-defined, thus making the resulting C*-algebras amenable to study. It had been known that, unlike to the group case, taking plain isometric representations leads to intractable objects, e.g., the universal isometric C*-algebra of $\mathbb{N}\times\mathbb{N}$ is not even nuclear [44]. A richer structure is at hand when one considers the relations in the Fock representation of a quasi-lattice order, now commonly known as Nica covariance. Nica’s setup is now covered by the theory as a product system with “trivial” fibers. On the other hand, Pimsner [49] introduced C*-algebras of product systems over $\mathbb{Z}_{+}$, i.e., C*-correspondences, that encodes successfully graph C*-algebras, crossed products over $\mathbb{Z}$, or partial transformations, see for example [32]. Motivated by [46, 49] as well as his work with Raeburn [21], Fowler [20] proceeded to an in-depth study of product systems over quasi-lattices, that inspired a great number of subsequent works. In order to make full use of this additional feature, Fowler imposed several axioms on the product system resulting to a by-default Wick ordering. Moving beyond the quasi-lattice order structure, Brownlowe, Larsen and Stammeier [7] considered product systems related to dynamical systems over right LCM semigroups. Kwaśniewski and Larsen [36, 37] extended this setup to general product systems over right LCM semigroups, and essentially this is the furthest one may go and still have a Wick ordering. Cuntz, Deninger and Laca [9] used the Fock model beyond Nica’s work [46] for $P=R\rtimes R^{\times}$, where Nica covariance may fail, but one can still use the Fock representation as the prototype for covariant relations.
@@
The coaction of the ambient group $G$ has been pivotal in the theory of product systems, in particular in relation to boundary quotients by Carlsen, Larsen, Sims and Vittadello [8] and Sehnem [52, 53], as well as in semigroup algebras by Laca and Sehnem [38] and the first author, Katsoulis, Laca and Li [25]. Dor-On, the first author, Katsoulis, Laca and Li [13] noted that there is a canonical Fell bundle arising from the normal coaction on ${\mathcal{T}}_{\lambda}(X)$, denoted here by ${\mathcal{F}}{\mathcal{C}}X$. It appears that, while ${\mathcal{T}}_{\lambda}(X)$ is the reduced C*-algebra of ${\mathcal{F}}{\mathcal{C}}X$, it is not always the reduced C*-algebra for the universal C*-algebra ${\mathcal{T}}(X)$ of the representations of $X$ (as it happens for $P=\mathbb{Z}_{+}$). To make a distinction, we will write ${\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$ for the universal C*-algebra of ${\mathcal{F}}{\mathcal{C}}X$; a representation of $X$ that promotes to a *-representation of ${\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$ will be called Fock covariant. It was shown in [13] that Fock covariant representations coincide with Nica covariant representations for a compactly aligned product system $X$ over a right LCM semigroup $P$, and thus ${\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$ coincides with the universal Nica covariant C*-algebra ${\mathcal{N}}{\mathcal{T}}(X)$ in this case.
@@
Motivated by the Hao–Ng isomorphism problem, our first aim in this work is to provide a description of the equivariant Fock covariant injective representations of $X$. For our purposes, we will view a product system $X$ as a family of C*-correspondences $\{X_{p}\}_{p\in P}$ in a common ${\mathcal{B}}(H)$ that satisfies some natural semigroup relations, i.e., that $A:=X_{e}$ is a C*-algebra, and
@@
This setup enables to still define left creation operators on the full Fock space ${\mathcal{F}}X$ giving rise to the Fock C*-algebra ${\mathcal{T}}_{\lambda}(X)$, while it recovers Fowler’s product systems in [20] which arise under the stronger saturation condition $[X_{p}\cdot X_{q}]=X_{pq}$ for all $p,q\in P$. We note that most of the works for product systems assume non-degeneracy of the left action on $X$. Nevertheless, non-degeneracy is rarely used, although imposed by the saturation condition of Fowler [20]. Our setting from (1.1) contains all non-degenerate product systems, while accommodating possible degenerate ones. Such cases arise naturally for example in C*-dynamical systems as the ones considered by Davidson, Fuller and the first named author in [10].
The characterisation of Fock covariance that we provide is given in terms of the $\mathbf{K}_{{\mathbf{x}}}$-cores for ${\mathbf{x}}\in{\mathcal{J}}$. We then use the natural description of the fixed point algebra by an inductive limit of C*-algebras on the $\cap$-closed finite subsets of the constructible ideals of $P$ in ${\mathcal{J}}$, which are spanned by the $\mathbf{K}_{\bullet}$-cores.
@@
We square this characterisation with established results from the literature. First, we provide a direct connection with the equivariant Nica covariant injective representations when $P$ is a right LCM semigroup which recovers [13, Proposition 4.3]. Secondly, when each $X_{p}=\mathbb{C}$ then we are at the setting of the semigroup representations and our description gives back the relations identified by Laca and Sehnem in [38, Section 3]; see also [26, Remark 3.15] (note that injectivity is automatic in the latter case). The connection with [13] and [38] can be seen categorically, however we give independent proofs from first principles that validate the compatibility with our results herein.
@@ 1.2. The Hao–Ng isomorphism problem
and then the C*-envelope theory of Katsoulis and Ramsey [30] can be implemented. In the absence of independence, it is unclear whether the *-isomorphism (1.5) still holds, but it does not exclude the possibility that the completely isometric isomorphism (1.6) is valid. A careful investigation of the arguments of [13, 14, 27, 28, 31] reveals that Fock covariance of the identity representation $X\rtimes_{\alpha,\lambda}{\mathfrak{H}}\hookrightarrow{\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ enables the completely isometric isomorphism (1.6). Our characterisation applies towards this resolution when ${\mathfrak{H}}$ is discrete without restrictions on the product system $X$ and the semigroup $P$. This recovers the aforementioned results appearing in [13, Theorem 6.3], [14, Theorem 6.6], and [27, Theorem 3.2].
@@ Acknowledgements.
The authors acknowledge that this research work was supported within the framework of the National Recovery and Resilience Plan Greece 2.0, funded by the European Union - NextGenerationEU (Implementation Body: HFRI. Project name: Noncommutative Analysis: Operator Systems and Nonlocality. HFRI Project Number: 015825). The second author acknowledges that this research work was supported by the Hellenic Foundation for Research and Innovation (HFRI) under the 5th Call for HFRI PhD Fellowships (Fellowship Number: 19145). The second author acknowledges that this publication is based upon work from COST Action CaLISTA CA21109 supported by COST (European Cooperation in Science and Technology), www.cost.eu. This manuscript will form part of the second author’s PhD thesis.
The authors would like to thank Elias Katsoulis for his helpful comments and remarks on the draft of the manuscript. The authors would like to thank the referee for the comments and remarks that helped improve the presentation.
@@ 2.1. Coactions on C*-algebras
There is a direct connection between coactions on C*-algebras, gradings and Fell bundles. The reader is addressed to [18, 19, 50] where this theory is laid in full detail. Since the group $G$ is discrete, the coactions we consider are automatically non-degenerate in the sense of [50], see for example [13, Remark 3.2].
@@
together spanning a dense subset of ${\mathcal{C}}$, see the proof of [45, Proposition 2.6]. If, in addition, the map $({\operatorname{id}}\otimes\lambda)\circ\delta$ is faithful, then $\delta$ will be called normal. It follows that $\delta$ is normal if and only if ${\mathcal{C}}$ admits a reduced coaction by $G$, i.e., there is a faithful *-homomorphism $\delta_{\lambda}\colon{\mathcal{C}}\to{\mathcal{C}}\otimes\mathrm{C}^{*}_{\lambda}(G)$ such that
@@
If there exists a conditional expectation $E\colon{\mathcal{C}}\to{\mathcal{C}}_{e}$, then the subspaces are linearly independent, see [18, Theorem 3.3]. In this case the C*-grading is called topological. By definition, a coaction on a C*-algebra induces a topological C*-grading.
@@
for the canonical embedding. We use the same symbol $\Psi$ for the *-homomorphism of $\mathrm{C}^{*}_{\max}({\mathcal{B}})$ induced by a representation $\{\Psi_{g}\}_{g\in G}$ of ${\mathcal{B}}$ (as $\Psi\circ\hat{j}_{g}=\Psi_{g}$). By universality, we have that $\{\hat{j}_{g}\}_{g\in G}$ is an equivariant representation of ${\mathcal{B}}$, and in particular by [19, Proposition 17.9] we have that the map $\hat{j}_{g}\colon{\mathcal{B}}_{g}\to[\mathrm{C}^{*}_{\max}({\mathcal{B}})]_{g}$ is an isometric isomorphism. Hence, any Fell bundle arises as a C*-grading from some C*-algebra.
@@
and thus a normal coaction. By [18, Theorem 3.3] we have that, if $\Psi$ is an equivariant representation of ${\mathcal{B}}$ that is injective on ${\mathcal{B}}_{e}$, then there are equivariant canonical *-epimorphisms
@@
More generally, let $\Psi$ be an equivariant (possibly non-injective) representation of ${\mathcal{B}}$, then by the proof of [19, Proposition 21.4] (or by using the Fourier transform) we obtain that every $\Psi_{g}$ has closed range, and hence we have the induced Fell bundle
@@
commutative, see [19, Proposition 21.2, Proposition 21.3]. We will make use of the following folklore fact for Fell bundles.
@@
It is obvious that, if the inclusion holds for all $g\in G$, then in particular it holds for $g=e$. Conversely, suppose that
@@
and so $x\in\ker\Psi_{2}$. Consequently we derive that $x\in\ker\Psi_{2}\cap[\mathrm{C}^{*}_{\max}({\mathcal{B}})]_{g}$, as required.
@@
In that case, the canonical quotient map $q_{\mathcal{I}}$ is equivariant, i.e., the coaction $\delta$ descends to a coaction on ${\mathcal{C}}/{\mathcal{I}}$, see [18, Proposition 3.11] and [8, Proposition A.1] for the full details. The following proposition is perhaps folklore, and we include a proof for completeness.
@@
On the other hand, since $q_{\mathcal{I}}({\mathcal{B}})$ is a topological C*-grading of $\mathrm{C}^{*}_{\max}({\mathcal{B}})/{\mathcal{I}}$, by [19, Theorem 19.5] there exists a canonical *-epimorphism
@@ 2.2. Semigroups
We will require some elements from the theory of semigroup algebras and right ideals of a semigroup from the work of Li [40], while we also fix notation. For a unital discrete left-cancellative semigroup $P$, we let the left creation (isometric) operators given by
@@
By definition we have $p^{-1}P=P$. We write ${\mathcal{J}}$ for the smallest family of right ideals of $P$ containing $P$ and $\emptyset$ that is closed under left multiplication, taking pre-images under left multiplication (as in the sense above) and finite intersections, i.e.,
@@
The right ideals of $P$ in ${\mathcal{J}}$ are called constructible. By the proof of [40, Lemma 3.3] we obtain a reduced form for the elements of ${\mathcal{J}}$, since
@@
Henceforth we will assume that $P$ is a unital subsemigroup of a discrete group $G$. In this case we have
@@
and therefore inductively we obtain
@@
is a reduced coaction, and thus it lifts to a normal coaction on $\mathrm{C}^{*}_{\lambda}(P)$. The induced faithful conditional expectation on $\mathrm{C}^{*}_{\lambda}(P)$ coincides with the compression to the diagonal $\ell^{\infty}(P)$. In the presence of an overlying group there is an explicit formula for the projection $E_{[{\mathbf{x}}]}$ for ${\mathbf{x}}\in{\mathcal{J}}$, obtained in [40, Lemma 3.1], i.e.,
@@ 2.3. C*-correspondences
The theory of Hilbert modules over C*-algebras is well-developed. The reader is addressed to [39, 43] for an excellent introduction to the subject; see also [32] and the references therein.